The Mars company says, that before the introduction of purple, yellow made up 20% of their plain M&M candies, red made up another 20%, and orange, blue, and green each made up 10%. The rest were brown. Assuming you had an infinite supply of M&M's with the older color distribution, if you picked six M&M's in a row, what is the probability that at least one is yellow?
Answer #1

P(Not a yellow candy) = 1 - 0.20 = 0.80

Hence,

P(Atleast one yellow out of 6)

= 1 - P(No yellow)

= 1 - (0.80)6

= 0.737856